# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v1

@@ Abstract
Band insulators can be classified into different topological phases according to different topological invariants [1, 2, 3, 4]. One important and basic topological invariant is the Chern number for two dimensional (2D) systems [2, 5, 1], which describes the winding number of hybrid Wannier centers of Bloch energy bands within 2D Brillouin zone [6], or the effective number of gapless chiral edge states [7, 6]. Those time-reversal-broken band insulators with nonzero Chern number are called Chern insulators, or quantum anomalous Hall (QAH) insulators, since they would exhibit quantum Hall effect in the absence of macroscopic magnetic field [8, 9, 10]. One can generalize the notion of 2D Chern insulators to its 3D counterpart by stacking an infinite number of 2D Chern insulators along the vertical axis [11, 12, 13, 14]. In principle, one can stack 2D Chern insulators along any direction so that one can define three independent Chern numbers $(C_{x},C_{y},C_{z})$, called Chern vector [15, 16], to characterize 3D band insulators. If the interlayer hopping is weak enough such that the bulk gap remains opened, such a 3D band insulator, named 3D Chern insulator, is adiabatically connected to an infinite number of decoupled 2D Chern-insulator layers.
@@
Topological 3D band insulators are known to exhibit non-trivial orbital magnetoelectric (OME) responses. The OME coupling $\alpha_{\mu\nu}$ is defined as the change in bulk electric polarization $P_{\mu}$ while applying magnetic field $B_{\nu}$, or equivalently the variation of bulk orbital magnetization $M_{\nu}$ induced by external electric field $\mathcal{E}_{\mu}$: $\alpha_{\mu\nu}=\partial P_{\mu}/\partial B_{\nu}=\partial M_{\nu}/\partial\mathcal{E}_{\mu}$ with $\mu,\nu=x,y,z$. This response is odd under inversion and time-reversal symmetry so that it vanishes if either of two symmetries is present. 3D time-reversal-protected topological insulators are classified by $\mathbb{Z}_{2}$ invariant [4, 17, 18, 19], which exhibits a half quantized bulk diagonal Chern-Simons OME response $\alpha_{\mu\nu}=e^{2}/2h\,\delta_{\mu\nu}$, modulo an integer multiple of $e^{2}/h$, once the surface state is gapped out [20, 21]. However, the Chern-Simons coupling is well defined only for bulk band insulators with zero Chern number within any 2D plane in momentum space, namely with vanishing Chern vector. This is because the expression of the Chern-Simons OME coupling coefficient involves an integration of Chern-Simons 3-form defined by the occupied Bloch states over the 3D Brillouin zone [21, 20], which requires the existence of a smooth and periodic gauge throughout the 3D Brillouin zone [22, 23, 24]. However, such a gauge is non-existing for 3D Chern insulators due to topological obstruction.
@@
Nevertheless, nothing forbids 3D Chern insulators to have a topological OME response. Intuitively, let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. What would happen if a non-disruptive electric field is applied assuming the interlayer coupling is weak? On the one hand, since 3D Chern insulator is adiabatically connected to trivial stacking of 2D Chern insulators, the off-diagonal $\alpha_{xz}$ or similar components should behave similarly as in the 2D case [25]. If a uniform electric field $\mathcal{E}_{x}$ is applied along the $x$ axis, a transverse anomalous Hall current flowing along the $y$ direction would be induced due to the non-zero Chern number, with $j_{y}=Ce^{2}/(dh)\mathcal{E}_{x}$. The bulk anomalous Hall current would naturally generate orbital magnetization $\mathbf{M}$ via ×$\nabla\times\mathbf{M}=\mathbf{j}$, giving rise to an off-diagonal OME response, with its response coefficient varying linearly in real space, i.e., $\partial_{x}\alpha_{zx}=-Ce^{2}/hd$ [25]. On the other hand, if a uniform electric field along the $z$ direction ($\mathcal{E}_{z}$) is applied, the surface-state occupation will be different for every layer due to the vertical electrostatic potential drop. Then, the uneven distribution of electrons among the chiral edge states of each Chern-insulator layer (in the weakly coupled situation) would cause a layer-resolved $M_{z}$, which may be manifested as a new type of diagonal OME effect which is unique to 3D Chern insulators, as schematically shown in Fig. 1(a).
@@
Figure 1: (a) Sketch map for quasi-3D Chern insulator. When $\mathcal{E}_{z}$ is applied, different layers have layer-dependent responses that the layer-resolved $M_{z}$s change linearly. (b) Honeycomb lattice for 2D Haldane model on which the hopping $t_{1}$ and $t_{2}\exp(i\phi)$ are marked. (c) Two ways of stacking: AA-stacking and AB-stacking. Unless mentioned ad hoc, we always use $t_{1}=1$, $t_{2}=1/3$, $\phi=\pi/3$, $t_{3}=0.2$, $t_{4}=0.15$ and the staggered mass term is set to zero.
@@
Although the total diagonal $\alpha_{zz}$, only relevant in 3D case, is vanishing if a mirror plane $\mathcal{M}_{z}$ exists, the layer-resolved $\alpha_{zz}^{(l)}$ (with layer index $l$), which measures the change of layer-resolved orbital magnetization $M_{z}^{(l)}$ under a weak $\mathcal{E}_{z}$, may be finite. To define $\alpha_{zz}^{(l)}$ in the thermodynamic limit, we need to define a layer-projected orbital magnetization $M_{z}^{(l)}$ using [26]
@@
where $\hat{P}=\sum_{E_{n}<\mu}|\psi_{n}\rangle\langle\psi_{n}|$ and $\hat{Q}=1-\hat{P}$ are the projection operators onto the occupied and unoccupied subspaces, respectively. The chemical potential $\mu$ is at charge neutrality, $x,y$ are 2D position operators and $A$ is the system area in the $x$-$y$ plane. “Tr” in the above equation means taking the trace over the whole Hilbert space. Since $\mathcal{E}_{z}$ would break the translation symmetry in the $z$ direction, it is natural to consider a slab geometry so that $M_{z}^{(l)}$ can be suitably expressed in the basis of the Bloch function [27]
@@
where $\rm o/e$ for occupied/empty states, $|u_{n\textbf{k}}\rangle$ is the periodic part of $|\psi_{n\textbf{k}}\rangle$ and $\epsilon_{z\alpha\beta}$ with $\alpha,\beta=x,y$ is the antisymmetric tensor. The layer-resolved magnetization of the $l$-th layer $M_{z}^{(l)}$ is then defined by summing over all the degrees of freedom except the layer index. For illustration, we model the 2D Chern-layer constituents of 3D Chern insulator by 2D Haldane model. Without loss of generality, the Chern number for each layer alone is set to be $C=-1$ and interlayer coupling is weak such that the bulk gap remains finite. Details of parameters is shown in Fig. 1(b). In our numerical calculations, we consider two types of stacking, $AA$ and $AB$, as illustrated in Fig. 1(c) and the slab consists of five or six layers. The k mesh is set to $100\times 100$. The chemical potential $\mu$ is determined by looking for charge neutrality point in the same system with open boundary conditions. Then, $\alpha_{zz}^{(l)}$ can be obtained numerically by finite difference, namely the ratio between change in $M_{z}^{(l)}$ and finite external electric field $\mathcal{E}_{z}$.
We find that $M_{z}^{(l)}$ varies linearly for weak electric field for both types of stacking, independent of number of layers (see Fig. 1 in Supplemental Materials [25]). The obtained $\alpha_{zz}^{(l)}$ is quantized to integers if $N_{\rm L}$ odd or to half-integers if $N_{\rm L}$ even, as shown in Table 1. The distribution is symmetric to the central plane of the system and the total $\alpha_{zz}$ is indeed vanishing as dictated by mirror symmetry. The quantization is only limited to $1\%$ precision due to the finite-size effect [25].
@@
Note that applying a vertical electric field induces a constant electrostatic potential drop between the two adjacent layers. Each layer acquires a layer-dependent on-site energy $e\mathcal{E}_{z}dl$. If $\mathcal{E}_{z}$ is so weak that the electrostatic potential drop between the top and bottom layers can be treated as a perturbation to the system, then $\alpha_{zz}^{(l)}$ can be determined semi-analytically using perturbation theory. We need to factorize all the changes linear in $\mathcal{E}_{z}$ among $|u_{n\textbf{k}}\rangle$, $C_{sl,n}(\textbf{k})$, $E_{n\textbf{k}}$ and $|\partial_{k_{\alpha}}u_{n\textbf{k}}\rangle$ and chemical potential $\mu$. Then, after considerable algebra, a semi-analytic formula including all the above four contributions for $\alpha_{zz}^{(l)}$ is derived. Details are given in Sec. III  of Supplemental Materials [25]. We say “semi-analytic” since the determination of the chemical potential still requires solving numerically the system with open boundary condition, which is unavoidable due to the presence of gapless surface states in 3D Chern insulators.
Therefore, one can directly calculate $\alpha_{zz}^{(l)}$ using only the energy and Bloch functions of the unperturbed system. As shown in the parenthesis of Table 1, $\alpha_{zz}^{(l)}$ obtained by semi-analytic perturbation theory follow exactly the same quantization rule as those calculated by numerical finite-difference method. Since the layer index $l$, the choice of which is a priori arbitrary, appears in perturbation formula [25], the final formula seems to depend on the “layer index gauge choice” at the first glance. Eventually, we can prove that the layer-resolved responses turns out to be “layer index gauge independent” as proved in Sec. IVof Supplemental Materials [25].
It worth a while to think that in a slab geometry, since the more layers there are, the greater the response coefficients on both sides, would not the corresponding $\alpha_{zz}^{(l)}$ diverge if the total number of layers went to infinity? However, when the total number of layers increases to a certain extent, even a very small electric field applied in the $z$ direction will cause Zener tunneling and the system then is no longer an insulator. This imposes a truncation to the otherwise diverging layer-resolved OME response.
@@
To understand the origin of the quantization in $\alpha_{zz}^{(l)}$, one can consider the simplest case where neighboring layers are completely decoupled from each other such that the Hamiltonian is block diagonal in layer space. The only contribution to $\alpha_{zz}^{(l)}$ comes from $E_{n\textbf{k}}-\mu$. Then, we can prove exactly that the difference in $\alpha_{zz}^{(l)}$ between the two neighboring layers is quantized (see Sec. V  of Supplemental Materials [25])
$\alpha_{zz}^{(l+1)}-\alpha_{zz}^{(l)}=C\,\frac{e^{2}}{h}\,,$ (4)
@@
$\frac{\partial\alpha_{zz}(l)}{\partial l}=C\,\frac{e^{2}}{h}\,.$ (5)
This quantization rule remains robust even in the presence of considerable interlayer hopping, as shown in Table 2 with $C=-1$. Knowing that the total $\alpha_{zz}$ is vanishing, we can deduce that $\alpha_{zz}^{(l)}$ is integer quantized in units of $e^{2}/h$ for odd number of layers or half-integer quantized for even number of layers, consistent with the previous findings. This is expected since the interlayer decoupled case is adiabatically connected to the above weakly coupled case.
The topological origin of the quantization rule is more transparent if we invoke the interpretation of OME coupling by the bulk-edge correspondence. It is precisely the chiral edge states that give rise to the quantized $\partial_{l}\alpha_{zz}(l)$ in a 3D Chern insulator. Intuitively, 3D Chern insulator is formed by stacking $N_{\rm L}$ 2D Chern insulators with, for example, $C=1$. There must be $N_{\rm L}$ chiral edge states flowing around each layer and a charge neutral $\mu$ crosses with all the edge states. A weak uniform $\mathcal{E}_{z}$ induces different occupation for the edge states of different layer when the system is back to equilibrium. The redistribution of electrons among different layers can be equivalently modelled by layer-resolved chemical potential $\mu^{(l)}=\mu_{0}+e\mathcal{E}_{z}d_{0}(l-\Delta l_{N_{\rm L}})$, where $\mu_{0}$ is the chemical potential in the absence of electric field and $\Delta l_{N_{\rm L}}=(N_{\rm L}+1)/2$ such that the total number of electrons is unchanged, namely $\sum\mu^{(l)}=N_{\rm L}\mu_{0}$. Since the chemical potential is in the bulk gap, all the changes in $M_{z}^{(l)}$ stems from the surface-state contribution [6, 27, 28, 29]:
$M_{z}^{(l)}=\mu^{(l)}\frac{2\pi e}{\hbar}C_{z}(l)\,.$ (6)
Therefore, we can get a simple formula for $\alpha_{zz}^{(l)}$ as follows
$\alpha_{zz}^{(l)}=\frac{\partial M_{z}^{(l)}}{\partial\mu^{(l)}}\frac{\partial\mu^{(l)}}{\partial\mathcal{E}}=C_{z}(l)\frac{e^{2}}{h}(l-\Delta l_{N_{\rm L}}),$ (7)
consistent with Eq. (4) knowing that $C_{z}(l)=C$. The behavior for even and odd number of layers are given precisely by the shift $\Delta l_{N_{\rm L}}$. This picture is exact in the interlayer decoupled case and can be generalizable to the weakly coupled case by the topological argument, as long as the bulk gap is not closed so that the layer Chern number remains quantized $C_{z}(l)=C$ [30, 31, 32, 33].
@@
The same quantization rules also exist for a genuine 3D Chern insulator in which the vertical stacking of 2D Chern layers goes to infinity with periodic boundary condition. To evaluate $\alpha_{zz}^{(l)}$, it is convenient to calculate the change of layer-resolved polarization $P_{z}^{(l)}$ [34] under a finite weak magnetic field $B_{z}$. By adopting a magnetic field commensurate with crystalline lattice, we restore the translational symmetry in all the three directions such that $\alpha_{zz}^{(l)}$ is determined in the thermodynamic limit without explicitly invoking finite-size effect in the slab geometry.
@@
where $\Omega$ is the size of the 2D magnetic unit cell, interlayer distance is constant $d=1$ and the averaged maximally localized Wannier center is proportional to the averaged Berry phase $\theta_{l}$ calculated by parallel-transport method [34, 35]
@@
where $N_{k}$ is the total number of k points in the $x$-$y$ plane. In practice, parallel-transport method gives $N_{\text{L}}N_{k}$ Berry phases, which are then categorized into $\theta_{l}$ of different layers based on the smallest distance criterion. Here $N_{\rm{L}}$ is the number of layers in an artificially chosen slab, and here we consider infinite stacking of slabs with periodic boundary condition imposed, so that the vertical size of 3D Brillouin zone is set to $2\pi/(N_{\rm{L}}d)$.
Using the same parameters as before, it turns out that even with a relative strong interlayer coupling ($t_{3},t_{4}$ is of the same order magnitude as $t_{2}$), the Wannier centers sit closely to each layer [25]. In practice, we use a magnetic field such that the magnetic flux piercing the atomic unit cell is $p/q=1/30$. A $48\times 48\times 48$ k mesh is adopted for numerical calculations. We obtain a similar quantization rule for $\alpha_{zz}^{(l)}$ as shown in Table 3. Compared with previous calculations, the quantization is way better ($\sim 0.1\%$) in such situation with fully periodic boundary condition.
@@
By Streda formula [36], when a finite magnetic field is turned on, states would flow between valence and conduction bands while keeping the chemical potential in the bulk gap. The number of transferred states is proportional to the Chern number of the system. For example, for $N_{\text{L}}$ = 4, when $B_{z}$ is switched on, four states would flow from conduction to valence bands as dictated by four $C=-1$ 2D Chern layers. Most saliently, we observe that these four states are tied exactly to four equidistant Wannier centers from the four coupled 2D Chern layers, respectively. Indeed, the quantization rule remains robust as long as the extra $B_{z}$-induced flowing states are evenly distributed among the $N_{\rm{L}}$ equidistant Wannier centers, which do not necessarily coincide with the layer position. The validity of Eq. (4) even do not require uniform layer charge density when $B_{z}=0$. The only requirement is that the bulk gap remains opened and layer Chern number remains exactly quantized as interlayer coupling is turned on, which is obvious for disorder-free 3D Chern insulators where $k_{z}$ is a good quantum number.
Following the above argument, here we give a more rigorous proof of the quantization rule Eq. (4) for interlayer-coupled 3D Chern insulators. Under finite $B_{z}$, due to Streda formula, the layer charge density $n_{l}$ follows $\partial n_{l}/\partial B_{z}=C_{z}(l)e^{2}/h$ for all layers, with $C_{z}(l)$ denoting layer Chern number. Therefore, whenever there is a change in the total charge density induced by $B_{z}$, it has to be equally distributed to each layer as long as $C_{z}(l)$ is still well defined and quantized. The additional layer charge density leads to a change in layer-resolved polarization $\delta P_{z}^{(l)}=\delta n_{l}\overline{z}_{l}/d$, which is induced by $B_{z}$. Thus, the layer-resolved OME coupling is given by:
$\alpha_{zz}^{(l)}=\frac{\delta\!P_{z}^{(l)}}{\delta\!n_{l}}\frac{\delta\!n_{l}}{\delta\!B_{z}}=\frac{\overline{z}_{l}}{d}C_{z}(l)\frac{e^{2}}{h}\,.$ (10)
Note that $\overline{z}_{l}$ does not necessarily sit at the corresponding layer position due to lack of symmetry. Let $\overline{z}_{l}=ld+\Delta z$ as in our cases with constant interlayer distance $d$, then Eq. (4) follows immediately. For illustration, let us consider a slab breaking inversion and mirror symmetry by introducing a different on-site energy only on the first layer (denoted as substrate layer) such that the total response $\alpha_{zz}$ is finite. In practice, the slab consists of five $AA$-stacked layers. We set the on-site energy of the first layer to be $1/6$ varying the amplitude of $t_{3,4}$. We calculate $\alpha_{zz}^{(l)}$ using Eq. (3) with $90\times 90$ $\mathbf{k}$ mesh. As shown in Table 4, although $\alpha_{zz}^{(l)}$ itself does not follow any quantization rule [25], the difference between the neighboring layer-resolved responses is quantized to layer Chern number as predicted. This shows that the most intrinsic quantization rule is Eq. (4). Nevertheless, symmetry can impose additional constraints on the values of $\Delta z$. For example, mirror or inversion symmetry forces $\Delta z$ to be integer or half integer, as observed in Table 1 and 3.
@@
In the presence of disorder, as long as the bulk gap is not closed by disorder such that layer Chern number remains well defined and quantized, the change of layer charge density induced by magnetic field would still remain quantized as dictated by Streda formula. Thus, the quantization rule of Eq. (4) is still valid. More numerical evidence of the robustness of Eq. (4) in the presence of disorder is given in Table  5. In practice, we need to calculate $\alpha_{zz}^{(l)}$ numerically by finite-difference in the fully open boundary conditions. Disorder is modelled by random on-site potential. Two types of disorder are introduced. The first type is that on-site energy only fluctuates in-plane repeating itself among all the layers. The second type breaks the translation symmetry such that on-site energy fluctuates among all the sites of all layers. We consider 15 disorder configurations and then analyze the results statistically. Here we set $t_{3}=0.2$, $t_{4}=0$, and a strong disorder with amplitude $0.4t_{3}$.
@@
Type-I fluctuation 2.73E-04 1.47E-04 1.34E-05 1.37E-04 2.82E-04

### v2

@@ Abstract
Band insulators can be classified into different topological phases according to different topological invariants [1, 2, 3, 4]. One important and basic topological invariant is the Chern number for two dimensional (2D) systems [2, 5, 1], which describes the winding number of hybrid Wannier centers within 2D Brillouin zone [6, 7].Those time-reversal-broken band insulators with nonzero Chern number are called Chern insulators, which would exhibit quantum anomalous Hall effect [8, 9, 10]. The notion of 2D Chern insulators can be generalized to its 3D counterpart by stacking an infinite number of 2D Chern insulators along the vertical axis [11, 12, 13, 14, 15]. In principle, one can stack 2D Chern insulators along any direction so that one can define three independent Chern numbers $(C_{x},C_{y},C_{z})$, called Chern vector [16, 17], to characterize 3D band insulators. If the interlayer hopping is weak enough such that the bulk gap remains opened, such a 3D band insulator, named 3D Chern insulator, is adiabatically connected to an infinite number of decoupled 2D Chern-insulator layers.
@@
Topological 3D band insulators are known to exhibit non-trivial orbital magnetoelectric (OME) responses. The OME coupling $\alpha_{\mu\nu}$ is defined as the change in bulk electric polarization $P_{\mu}$ while applying magnetic field $B_{\nu}$, or equivalently the variation of bulk orbital magnetization $M_{\nu}$ induced by external electric field $\mathcal{E}_{\mu}$: $\alpha_{\mu\nu}=\partial P_{\mu}/\partial B_{\nu}=\partial M_{\nu}/\partial\mathcal{E}_{\mu}$ with $\mu,\nu=x,y,z$. This response is odd under inversion and time-reversal symmetry so that it vanishes if either of two symmetries is present. 3D time-reversal-protected topological insulators are classified by $\mathbb{Z}_{2}$ invariant [4, 18, 19, 20], which exhibits a half quantized bulk diagonal Chern-Simons OME response $\alpha_{\mu\nu}=e^{2}/2h\,\delta_{\mu\nu}$, modulo an integer multiple of $e^{2}/h$, once the surface state is gapped out [21, 22]. However, the Chern-Simons coupling is well defined only for bulk band insulators with zero Chern number within any 2D plane in momentum space, namely with vanishing Chern vector. This is because the expression of the Chern-Simons OME coupling coefficient involves an integration of Chern-Simons 3-form defined by the occupied Bloch states over the 3D Brillouin zone [22, 21, 23, 24, 25], which requires the existence of a smooth and periodic gauge throughout the 3D Brillouin zone [26, 27, 28]. However, such a gauge is non-existing for 3D Chern insulators due to topological obstruction.
@@
Nevertheless, nothing forbids 3D Chern insulators to have a topological OME response. Intuitively, let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. On the one hand, it is straightforward to see that the quantized anomalous Hall current in response to weak in-plane electric field would induce to an off-diagonal OME response with its coefficient varying linearly in reap space (see Sec. VII in Supplemental Material). On the other hand, if an out-of-plane uniform electric field ($\mathcal{E}_{z}$) is applied, the surface-state occupation would be different for every layer due to the vertical electrostatic potential drop. Then, the uneven distribution of electrons among the chiral edge states of each weakly-coupled Chern-insulator layer would cause a layer-resolved orbital magnetization ($M_{z}$), which may be manifested as a new type of diagonal OME effect unique to 3D Chern insulators, as schematically shown in Fig. 1(a).
@@
Figure 1: (a) Schematic for quasi-3D Chern insulator (see text). (b) Honeycomb lattice for 2D Haldane model on which the hopping $t_{1}$ and $t_{2}\exp(i\phi)$ are marked. (c) Two ways of stacking: $AA$-stacking and $AB$-stacking. Unless mentioned ad hoc, we always use $t_{1}=1$, $t_{2}=1/3$, $\phi=\pi/3$, $t_{3}=0.2$, $t_{4}=0.15$ and the staggered mass term is zero.
@@
Although the total diagonal $\alpha_{zz}$, only relevant in 3D case, is vanishing in the presence of mirror or inversion symmetry, the layer-resolved $\alpha_{zz}^{(l)}$ (with layer index $l$), which measures the change of layer-resolved orbital magnetization $M_{z}^{(l)}$ under a weak $\mathcal{E}_{z}$, may be finite. To define $\alpha_{zz}^{(l)}$ in the thermodynamic limit, we need to define a layer-projected orbital magnetization $M_{z}^{(l)}$ using [29, 30]
@@
where $\hat{P}=\sum_{E_{n}<\mu}|\psi_{n}\rangle\langle\psi_{n}|$ and $\hat{Q}=1-\hat{P}$ are the projection operators onto the occupied and unoccupied subspaces, respectively. The chemical potential $\mu$ is at charge neutrality, $x,y$ are 2D position operators and $A$ is the system area in the $x$-$y$ plane. “Tr” in the above equation means taking the trace over the whole Hilbert space. Since $\mathcal{E}_{z}$ would break the translational symmetry in the $z$ direction, it is natural to consider a slab geometry so that $M_{z}^{(l)}$ can be suitably expressed in the basis of the Bloch function
@@
where $\rm o/e$ for occupied/empty states, $|u_{n\textbf{k}}\rangle$ is the periodic part of $|\psi_{n\textbf{k}}\rangle$ and $\epsilon_{z\alpha\beta}$ with $\alpha,\beta=x,y$ is the antisymmetric tensor. The layer-resolved magnetization of the $l$-th layer $M_{z}^{(l)}$ is then defined by summing over all the degrees of freedom except the layer index. For illustration, we model the 2D Chern-layer constituents of 3D Chern insulator by 2D Haldane model. Without loss of generality, the Chern number for each layer is set to be $C=1$ and interlayer coupling is weak such that the bulk gap remains finite. Details of parameters is shown in Fig. 1(b). In our numerical calculations, we consider two types of stacking, $AA$ and $AB$, as illustrated in Fig. 1(c) and the slab consists of five or six layers. The k mesh is set to $100\times 100$. The chemical potential $\mu$ is determined by looking for charge neutrality point in the same system with open boundary conditions. Then, $\alpha_{zz}^{(l)}$ can be obtained by numerically taking the ratio between change of $M_{z}^{(l)}$ and $\mathcal{E}_{z}$. We find that $M_{z}^{(l)}$ varies linearly for weak electric field for both types of stacking (see Fig. 1 in Supplemental Material [31]). The obtained $\alpha_{zz}^{(l)}$ is quantized to integers if $N_{\rm L}$ odd or to half-integers if $N_{\rm L}$ even, as shown in Table 2. The distribution is symmetric to the central plane of the system and the total $\alpha_{zz}$ is indeed vanishing as dictated by mirror symmetry. The quantization is limited to $1\%$ precision due to the finite-size effect [31].
@@
If $\mathcal{E}_{z}$ is so weak that the electrostatic potential drop between the top and bottom layers can be treated as a perturbation to the system, then $\alpha_{zz}^{(l)}$ can be determined semi-analytically using perturbation theory. After considerable algebra, a semi-analytic formula for $\alpha_{zz}^{(l)}$ is derived. Details are given in Sec. III  of Supplemental Material [31]. We say “semi-analytic” since the determination of the chemical potential still requires solving numerically the system with open boundary condition, which is unavoidable due to the presence of gapless surface states in 3D Chern insulators. Therefore, one can directly calculate $\alpha_{zz}^{(l)}$ using only the energy and Bloch functions of the unperturbed system. As shown in the parenthesis of Table 2, $\alpha_{zz}^{(l)}$ obtained by semi-analytic perturbation theory follow exactly the same quantization rule as those calculated by numerical finite-difference method.
@@
To understand the origin of the quantization in $\alpha_{zz}^{(l)}$, one can consider the simplest case where neighboring layers are completely decoupled from each other such that the Hamiltonian is block diagonal in layer space. The only contribution to $\alpha_{zz}^{(l)}$ comes from $E_{n\textbf{k}}-\mu$ (see Eq. (3)). Then, we can prove exactly that the difference in $\alpha_{zz}^{(l)}$ between the two neighboring layers is quantized (see Sec. V  of Supplemental Material [31])
$\alpha_{zz}^{(l+1)}-\alpha_{zz}^{(l)}=-C\,\frac{e^{2}}{h}\,,$ (4)
@@
$\frac{\partial\alpha_{zz}(l)}{\partial l}=-C\,\frac{e^{2}}{h}\,.$ (5)
This quantization rule remains robust even in the presence of considerable interlayer hopping, as shown in Table 2 with $C=1$. Knowing that the total $\alpha_{zz}$ is vanishing, we can deduce that $\alpha_{zz}^{(l)}$ is integer quantized in units of $e^{2}/h$ for odd number of layers or half-integer quantized for even number of layers, consistent with the previous findings.
It is the topological edge states that give rise to the quantized $\partial_{l}\alpha_{zz}(l)$ in a 3D Chern insulator. Intuitively, 3D Chern insulator is formed by stacking $N_{\rm L}$ 2D Chern insulators. There must be $N_{\rm L}$ gapless chiral edge states and a charge neutral $\mu$ crosses with all the edge states. A weak uniform $\mathcal{E}_{z}$ induces different occupation for the edge states of different layer when the system is back to equilibrium. The redistribution of electrons among different layers can be equivalently modelled by layer-resolved chemical potential $\mu^{(l)}=\mu_{0}-e\mathcal{E}_{z}d(l-\Delta l_{N_{\rm L}})$, where $\mu_{0}$ is the chemical potential in the absence of electric field and $\Delta l_{N_{\rm L}}=(N_{\rm L}+1)/2$ such that the total number of electrons is unchanged, namely $\sum\mu^{(l)}=N_{\rm L}\mu_{0}$. Since the chemical potential is in the bulk gap, all the changes in $M_{z}^{(l)}$ stems from the surface-state contribution [6, 30, 32, 33]:
$M_{z}^{(l)}=\mu^{(l)}\frac{e}{hd}C_{z}(l)\,.$ (6)
where the division by $d$ makes $M_{z}^{(l)}$ volumetric. Therefore, we can get a simple formula for $\alpha_{zz}^{(l)}$ as follows
$\alpha_{zz}^{(l)}=\frac{\partial M_{z}^{(l)}}{\partial\mu^{(l)}}\frac{\partial\mu^{(l)}}{\partial\mathcal{E}_{z}}=-C_{z}(l)\frac{e^{2}}{h}(l-\Delta l_{N_{\rm L}}),$ (7)
consistent with Eq. (4) knowing that $C_{z}(l)=C$. The behavior for even and odd number of layers are given precisely by the shift $\Delta l_{N_{\rm L}}$. This picture is exact in the interlayer decoupled case and generalizable to the interlayer coupled case by the topological argument, as long as the bulk gap is not closed so that the layer Chern number remains quantized $C_{z}(l)=C$ [34, 35, 36, 37].
@@
The same quantization rule also exists for a genuine 3D Chern insulator in which the vertical stacking of 2D Chern layers goes to infinity with periodic boundary condition. To evaluate $\alpha_{zz}^{(l)}$, it is convenient to calculate the change of layer-resolved polarization $P_{z}^{(l)}$ [38] under a finite weak magnetic field $B_{z}$. By adopting a magnetic field commensurate with crystalline lattice, we restore the translational symmetry in all the three directions such that $\alpha_{zz}^{(l)}$ is determined in the thermodynamic limit without explicitly invoking finite-size effect in the slab geometry.
@@
where $\Omega$ is the area of the 2D magnetic unit cell, interlayer distance is constant $d=1$ and the averaged maximally localized Wannier center is proportional to the averaged Berry phase $\theta_{l}$ calculated by parallel-transport method [38, 39]
@@
where $N_{k}$ is the total number of k points in the $x$-$y$ plane. In practice, parallel-transport method gives $N_{\text{L}}N_{k}$ Berry phases, which are then categorized into $\theta_{l}$ of different layers based on the smallest distance criterion. Here $N_{\rm{L}}$ is the number of layers in an artificially chosen slab, and we consider infinite stacking of slabs with periodic boundary condition imposed, so that the vertical size of 3D Brillouin zone is set to $2\pi/(N_{\rm{L}}d)$.
It turns out that even with a relative strong interlayer coupling ($t_{3},t_{4}$ is of the same order magnitude as $t_{2}$), the Wannier centers still sit closely to each layer [31]. In practice, we use a magnetic field such that the magnetic flux piercing the atomic unit cell is $p/q=1/30$. A $48\times 48\times 48$ k mesh is adopted for numerical calculations. We obtain a similar quantization rule for $\alpha_{zz}^{(l)}$ as shown in Table 3. Compared with previous calculations, the quantization is much better ($\sim 0.1\%$) in such situation with fully periodic boundary condition.
@@
Here we give a somewhat rigorous proof of the quantization rule Eq. (4) for interlayer-coupled 3D Chern insulators. Under finite $B_{z}$, due to Streda formula, the layer charge density $n_{l}$ follows $\partial n_{l}/\partial B_{z}=-C_{z}(l)e^{2}/h$ for all layers, with $C_{z}(l)$ denoting layer Chern number. Therefore, by Streda formula, whenever there is a change in the total charge density (of the $N_{\rm{L}}$-layer unit cell) induced by $B_{z}$, it has to be evenly distributed to the $N_{\rm{L}}$ equidistant averaged Wannier centers, as long as $C_{z}(l)$ is still well defined and quantized. The additional layer charge density leads to a change in layer-resolved polarization $\delta P_{z}^{(l)}=\delta n_{l}\overline{z}_{l}/d$ (according to Eq. (8)), which is induced by $B_{z}$. Thus, the layer-resolved OME coupling is given by:
$\alpha_{zz}^{(l)}=\frac{\delta P_{z}^{(l)}}{\delta n_{l}}\frac{\delta n_{l}}{\delta B_{z}}=-\frac{\overline{z}_{l}}{d}C_{z}(l)\frac{e^{2}}{h}\,.$ (10)
Note that the average Wannier center $\overline{z}_{l}$ does not necessarily sit at the corresponding layer position due to lack of symmetry. Let $\overline{z}_{l}=ld+\Delta z$ as in our cases with constant interlayer distance $d$, then Eq. (4) follows immediately. Actually the validity of Eq. (4) even does not require uniform layer charge density when $B_{z}=0$. The only requirement is that the bulk gap remains opened and layer Chern number remains exactly quantized as interlayer coupling is turned on, which is obvious for disorder-free 3D Chern insulators where $k_{z}$ is a good quantum number.
For illustration, let us consider a slab breaking inversion and mirror symmetries by introducing a different on-site energy only to the first layer (denoted as substrate layer) such that the total response $\alpha_{zz}$ becomes finite. In practice, the slab consists of five $AA$-stacked layers. We set the on-site energy of the first layer to be $1/6$ while varying the amplitude of $t_{3,4}$. We calculate $\alpha_{zz}^{(l)}$ using Eq. (3) with $90\times 90$ $\mathbf{k}$ mesh. As shown in Table 4, although $\alpha_{zz}^{(l)}$ itself does not follow any quantization rule [31], the difference between the neighboring layer-resolved responses is quantized to layer Chern number as predicted. This shows that the most intrinsic quantization rule is Eq. (4). Nevertheless, symmetry can impose additional constraints on the values of $\Delta z$. For example, mirror or inversion symmetry forces $\Delta z$ to be integer or half integer, as observed in Table 2 and 3.
@@
Type-I fluctuation 2.73×$\times 10^{-4}$ 1.47×$\times 10^{-4}$ 1.34×$\times 10^{-5}$ 1.37×$\times 10^{-4}$ 2.82×$\times 10^{-4}$